# Mind Meets Robots: A Review of EEG-Based Brain-Robot Interaction Systems

### v1

@@ 4.1.2. Signal Decoding
(Belkacem and Lakas, 2021; Chhabra et al., 2020; Wu et al., 2022; Sugiyama et al., 2023; Toichoa Eyam et al., 2021; Tang et al., 2022; Liu and Jebelli, 2021; Mondini et al., 2020; Xu et al., 2018; Korovesis et al., 2019; Karunasena et al., 2021; Chen et al., 2020; Wang et al., 2018b; Jo et al., 2022; Aldini et al., 2022; Rashid et al., 2020; Guo et al., 2020; Zhao et al., 2020b; Ehrlich and Cheng, 2019; Prinsen et al., 2022; Roshdy et al., 2021; Bahman and Shamsollahi, 2019; Francis et al., 2021)
High Accuracy
15 (17.2%)
@@
(Rahul and Sharma, 2019; Wang et al., 2018c; Cao and Liu, 2018; Martínez-Cagigal et al., 2020; Magee and Givigi, 2021; Chen et al., 2021; Roy and Bhaumik, 2022; Jo et al., 2022; Aldini et al., 2022; Rashid et al., 2020; Guo et al., 2020; Francis et al., 2021; Abougarair et al., 2021; Penaloza et al., 2018; Nann et al., 2020)
High Temporal Resolution
8 (9.2%)
@@
(Baka et al., 2019; Pawuś and Paszkiel, 2022; Aznan et al., 2019; Kim et al., 2021; Farmaki et al., 2022; Li et al., 2019; Zhao et al., 2020b; Ehrlich and Cheng, 2019)
Seamless Feedback
42 (48.3%)
@@
(Araujo et al., 2021; Chen et al., 2021; Roy and Bhaumik, 2022; Kim et al., 2021; Farmaki et al., 2022; Li et al., 2019; Prinsen et al., 2022; Roshdy et al., 2021; Bahman and Shamsollahi, 2019; Braun et al., 2019; Li et al., 2023; Qian et al., 2018; Dissanayake et al., 2022; Chang and Sun, 2021; Aldini et al., 2019; Staffa and Rossi, 2022; Iwane et al., 2019; Kompatsiari et al., 2018; Long et al., 2019; Yoon et al., 2021; Ogino and Mitsukura, 2018; Cervantes et al., 2023; Kar et al., 2021; Chu et al., 2018; Kilmarx et al., 2018; Li et al., 2022a; Lu et al., 2020; Zhang et al., 2021; Du et al., 2021; Sanguantrakul et al., 2020; Hernandez-Carmona and Penaloza, 2019; Chiuzbaian et al., 2019; Ghosh and Orlando, 2019; Wei et al., 2021; Ali et al., 2021; Gordleeva et al., 2020; Kuffuor and Samanta, 2018; Memar and Esfahani, 2018; Richter et al., 2023; Abougarair et al., 2021; Penaloza et al., 2018; Nann et al., 2020)
Computational Efficiency
4 (4.6%)
(Lu et al., 2022; Rahul and Sharma, 2019; Araujo et al., 2021; Richter et al., 2023)
Feedback Enhancement
Visual Feedback
28 (32.2%)
@@
(Baka et al., 2019; Belkacem and Lakas, 2021; Farmaki et al., 2022; Sugiyama et al., 2023; Kompatsiari et al., 2018; Si-Mohammed et al., 2018; Chenga et al., 2023; Kilmarx et al., 2018; Wang et al., 2018c; Cao and Liu, 2018; Abougarair et al., 2021; Du et al., 2021; Magee and Givigi, 2021; Penaloza et al., 2018; Yu et al., 2019; Chiuzbaian et al., 2019; Bahman and Shamsollahi, 2019; Ghosh and Orlando, 2019; Mondini et al., 2020; Rahul and Sharma, 2019; Yuan and Li, 2018; Quiles et al., 2022; Prinsen et al., 2022; Aznan et al., 2019; Hernandez-Carmona and Penaloza, 2019; Zhang et al., 2021; Ehrlich and Cheng, 2019; Alimardani et al., 2022)
Auditory Feedback
3 (3.4%)
(Kar et al., 2021; Richter et al., 2023; Wang and Sugaya, 2021)
Multimodal Feedback
7 (8%)
(Braun et al., 2019; Li et al., 2023; Chen et al., 2021; Li et al., 2022a; Roy and Bhaumik, 2022; Li et al., 2019; Dissanayake et al., 2022)
Direct Feedback
22 (25.3%)
@@
(Prinsen et al., 2022; Aznan et al., 2019; Hernandez-Carmona and Penaloza, 2019; Zhang et al., 2021; Ehrlich and Cheng, 2019; Alimardani et al., 2022; Chhabra et al., 2020; Fang et al., 2022; Memar and Esfahani, 2018; Lu et al., 2022; Pawuś and Paszkiel, 2022; Jo et al., 2022; Aldini et al., 2022; Martínez-Cagigal et al., 2020; Sanguantrakul et al., 2020; Wei et al., 2021; Araujo et al., 2021; Francis et al., 2021; Rashid et al., 2020; Staffa and Rossi, 2022; Cervantes et al., 2023; Kar et al., 2021)
Performance Feedback
18 (20.1%)
@@
(Cervantes et al., 2023; Guo et al., 2020; Qian et al., 2018; Wu et al., 2022; Iwane et al., 2019; Memar and Esfahani, 2019; Long et al., 2019; Liu and Jebelli, 2021; Lu et al., 2020; Shao et al., 2020; Boonarchatong and Ketcham, 2023; Kuffuor and Samanta, 2018; Jiang et al., 2018; Chen et al., 2020; Wang et al., 2018b; Ai et al., 2018; Lyu et al., 2022; Aldini et al., 2019)
Neurofeedback
22 (25.3%)
@@
(Staffa and Rossi, 2022; Kim et al., 2021; Dissanayake et al., 2022; Chang and Sun, 2021; Toichoa Eyam et al., 2021; Alimardani et al., 2020; Yoon et al., 2021; Tang et al., 2022; Ogino and Mitsukura, 2018; Chu et al., 2018; Roshdy et al., 2021; Ali et al., 2021; Xu et al., 2018; Ak et al., 2022; Zhao et al., 2020b; Gordleeva et al., 2020; Nann et al., 2020; Korovesis et al., 2019; Karunasena et al., 2021; Guo et al., 2020; Lyu et al., 2022; Aldini et al., 2019; Richter et al., 2023)
Artificial Intelligence Methods
Traditional Machine Learning (ML)
42 (48.3%)
@@
(Araujo et al., 2021; Cao and Liu, 2018; Magee and Givigi, 2021; Chen et al., 2021; Roy and Bhaumik, 2022; Farmaki et al., 2022; Li et al., 2019; Belkacem and Lakas, 2021; Chhabra et al., 2020; Xu et al., 2018; Chen et al., 2020; Wang et al., 2018b; Jo et al., 2022; Aldini et al., 2022; Rashid et al., 2020; Guo et al., 2020; Zhao et al., 2020b; Ehrlich and Cheng, 2019; Francis et al., 2021; Braun et al., 2019; Qian et al., 2018; Chang and Sun, 2021; Iwane et al., 2019; Long et al., 2019; Ogino and Mitsukura, 2018; Kar et al., 2021; Chu et al., 2018; Lu et al., 2020; Zhang et al., 2021; Du et al., 2021; Wei et al., 2021; Gordleeva et al., 2020; Kuffuor and Samanta, 2018; Penaloza et al., 2018; Lyu et al., 2022; Memar and Esfahani, 2019; Si-Mohammed et al., 2018; Shao et al., 2020; Boonarchatong and Ketcham, 2023; Yuan and Li, 2018; Jiang et al., 2018; Quiles et al., 2022)
Deep Learning (DL)
19 (21.8%)
@@
(Rahul and Sharma, 2019; Magee and Givigi, 2021; Pawuś and Paszkiel, 2022; Aznan et al., 2019; Li et al., 2019; Wu et al., 2022; Tang et al., 2022; Korovesis et al., 2019; Aldini et al., 2022; Guo et al., 2020; Roshdy et al., 2021; Li et al., 2023; Staffa and Rossi, 2022; Kar et al., 2021; Lu et al., 2020; Sanguantrakul et al., 2020; Ghosh and Orlando, 2019; Fang et al., 2022; Chenga et al., 2023)
Non-ML/DL
7 (8%)
@@
(Baka et al., 2019; Wang et al., 2018c; Yoon et al., 2021; Li et al., 2022a; Hernandez-Carmona and Penaloza, 2019; Chiuzbaian et al., 2019; Wang and Sugaya, 2021)
@@ 4.2. Robot
Context Category Domain Description Papers (%) References
Human Control Concern (HCC)
Human-centric Technology
Usability and user experience
Exploring better UX for effective interactions
7 (8%)
(Fang et al., 2022; Kim et al., 2021; Martínez-Cagigal et al., 2020; Kuffuor and Samanta, 2018; Yuan and Li, 2018; Memar and Esfahani, 2019; Iwane et al., 2019)
Research and innovation
Exploring new advancements for research needs
18 (20.7%)
(Ehrlich and Cheng, 2019; Sugiyama et al., 2023; Kilmarx et al., 2018; Cao and Liu, 2018; Chen et al., 2021; Liu and Jebelli, 2021; Chen et al., 2020; Zhao et al., 2020b; Hernandez-Carmona and Penaloza, 2019; Ak et al., 2022; Guo et al., 2020; Yuan and Li, 2018; Jiang et al., 2018; Si-Mohammed et al., 2018; Pawuś and Paszkiel, 2022; Wei et al., 2021; Cervantes et al., 2023; Lu et al., 2020)
Mutual collaboration
Collaborating with robots to enhance task performance by control
3 (3.4%)
(Lyu et al., 2022; Liu and Jebelli, 2021; Memar and Esfahani, 2018)
Assistance
Physical movement
Aiding people with motor disorders for physical movements
19 (21.8%)
(Farmaki et al., 2022; Zhang et al., 2021; Du et al., 2021; Abougarair et al., 2021; Magee and Givigi, 2021; Wang et al., 2018b; Li et al., 2019; Chiuzbaian et al., 2019; Ghosh and Orlando, 2019; Mondini et al., 2020; Rahul and Sharma, 2019; Ali et al., 2021; Francis et al., 2021; Bahman and Shamsollahi, 2019; Li et al., 2023; Long et al., 2019; Wang et al., 2018c; Tang et al., 2022; Wei et al., 2021)
Communication
Motionless control from people for intended actions
7 (8%)
(Boonarchatong and Ketcham, 2023; Shao et al., 2020; Rashid et al., 2020; Korovesis et al., 2019; Karunasena et al., 2021; Quiles et al., 2022; Bahman and Shamsollahi, 2019)
Healthcare
Rehabilitation
Rehabilitation for people with mobility impairment
15 (17.2%)
(Wu et al., 2022; Chu et al., 2018; Ai et al., 2018; Li et al., 2022a; Roy and Bhaumik, 2022; Jo et al., 2022; Nann et al., 2020; Xu et al., 2018; Sanguantrakul et al., 2020; Araujo et al., 2021; Gordleeva et al., 2020; Chhabra et al., 2020; Dissanayake et al., 2022; Wang et al., 2018c; Tang et al., 2022)
Therapy
Therapeutic process
1 (1.1%)
(Roshdy et al., 2021)
Entertainment
Engagement in entertaining activities
1 (1.1%)
(Cervantes et al., 2023)
Safety and Security
Exploration for safe and secure needs
1 (1.1%)
(Lu et al., 2020)
Social Interaction
Robots as companions for social communications
1 (1.1%)
(Toichoa Eyam et al., 2021)
Military Usage
Exploration intended for military cases
1 (1.1%)
(Belkacem and Lakas, 2021)
Robot Design Concern (RDC)
Human-centric Technology
Usability and user experience
Exploring better UX for effective interactions
2 (2.3%)
(Aldini et al., 2019; Chenga et al., 2023)
Research and innovation
Exploring new advancements for research needs
5 (5.7%)
(Richter et al., 2023; Penaloza et al., 2018; Lu et al., 2022; Kompatsiari et al., 2018; Ogino and Mitsukura, 2018)
Mutual collaboration
Collaborating with robots to enhance task performance
2 (2.3%)
(Aldini et al., 2022, 2019)
Assistance
Physical movement
Aiding people with motor disorders for physical movements
3 (3.4%)
(Yu et al., 2019; Aznan et al., 2019; Yoon et al., 2021)
Communication
Aiding motionless people for intended actions
2 (2.3%)
(Braun et al., 2019; Wang and Sugaya, 2021)
Healthcare
Rehabilitation
Rehabilitation for people with mobility impairment
1 (1.1%)
(Chenga et al., 2023)
Mental recovery
Supports mental recovery
1 (1.1%)
(Yoon et al., 2021)
Meditation
Help with mindfulness and meditation process
1 (1.1%)
(Yoon et al., 2021)
Education
Language learning
Supports exclusively in language learning
2 (2.3%)
(Prinsen et al., 2022; Alimardani et al., 2022)
Self-learning
Support self-learning for skill enrichment
3 (3.4%)
(Wang and Sugaya, 2021; Kar et al., 2021; Alimardani et al., 2020)
Entertainment
Engagement in entertaining activities
1 (1.1%)
(Kar et al., 2021)
Safety and Security
Exploration for safe and secure needs
1 (1.1%)
(Qian et al., 2018)
Social Interaction
Robots as companions for social communications
4 (4.6%)
(Alimardani et al., 2020; Baka et al., 2019; Chang and Sun, 2021; Staffa and Rossi, 2022)
@@ 4.3.1. Application Contexts
Table 5. The evaluation methods identified in our corpus.
Metrics Papers (%) References
Evaluation Methods
Performance metrics
54 (62.1%)
(Belkacem and Lakas, 2021; Memar and Esfahani, 2018; Staffa and Rossi, 2022; Iwane et al., 2019; Kompatsiari et al., 2018; Long et al., 2019; Si-Mohammed et al., 2018; Chenga et al., 2023; Tang et al., 2022; Pawuś and Paszkiel, 2022; Kar et al., 2021; Chu et al., 2018; Kilmarx et al., 2018; Cao and Liu, 2018; Chen et al., 2021; Jo et al., 2022; Aldini et al., 2022; Lu et al., 2020; Zhang et al., 2021; Du et al., 2021; Magee and Givigi, 2021; Roshdy et al., 2021; Aznan et al., 2019; Sanguantrakul et al., 2020; Penaloza et al., 2018; Yu et al., 2019; Bahman and Shamsollahi, 2019; Rashid et al., 2020; Zhao et al., 2020b; Guo et al., 2020; Nann et al., 2020; Gordleeva et al., 2020; Korovesis et al., 2019; Karunasena et al., 2021; Kuffuor and Samanta, 2018; Jiang et al., 2018; Chen et al., 2020; Wang et al., 2018b; Quiles et al., 2022; Mondini et al., 2020; Wei et al., 2021; Araujo et al., 2021; Rahul and Sharma, 2019; Francis et al., 2021; Shao et al., 2020; Xu et al., 2018; Boonarchatong and Ketcham, 2023; Prinsen et al., 2022; Li et al., 2023; Qian et al., 2018; Farmaki et al., 2022; Lyu et al., 2022; Wu et al., 2022; Sugiyama et al., 2023)
UX metrics
16 (18.4%)
(Baka et al., 2019; Memar and Esfahani, 2018; Chang and Sun, 2021; Toichoa Eyam et al., 2021; Staffa and Rossi, 2022; Alimardani et al., 2020; Chenga et al., 2023; Ogino and Mitsukura, 2018; Cervantes et al., 2023; Martínez-Cagigal et al., 2020; Nann et al., 2020; Li et al., 2023; Chhabra et al., 2020; Fang et al., 2022; Lyu et al., 2022; Richter et al., 2023)
Signal information
20 (23%)
(Ehrlich and Cheng, 2019; Dissanayake et al., 2022; Aldini et al., 2019; Wang and Sugaya, 2021; Lu et al., 2022; Alimardani et al., 2022; Chenga et al., 2023; Chu et al., 2018; Li et al., 2022a; Roy and Bhaumik, 2022; Aldini et al., 2022; Penaloza et al., 2018; Yu et al., 2019; Li et al., 2019; Chiuzbaian et al., 2019; Karunasena et al., 2021; Yuan and Li, 2018; Ai et al., 2018; Mondini et al., 2020; Kim et al., 2021)

### v3

@@ 4.1.2. Signal Decoding
Real-time Requirements Low Latency 23 (26.4%) (Belkacem and Lakas, 2021; Chhabra et al., 2020; Wu et al., 2022; Sugiyama et al., 2023; Toichoa Eyam et al., 2021; Tang et al., 2022; Liu and Jebelli, 2021; Mondini et al., 2020; Xu et al., 2018; Korovesis et al., 2019; Karunasena et al., 2021; Chen et al., 2020; Wang et al., 2018b; Jo et al., 2022; Aldini et al., 2022; Rashid et al., 2020; Guo et al., 2020; Zhao et al., 2020b; Ehrlich and Cheng, 2019; Prinsen et al., 2022; Roshdy et al., 2021; Bahman and Shamsollahi, 2019; Francis et al., 2021)
@@
High Accuracy 15 (17.2%) (Rahul and Sharma, 2019; Wang et al., 2018c; Cao and Liu, 2018; Martínez-Cagigal et al., 2020; Magee and Givigi, 2021; Chen et al., 2021; Roy and Bhaumik, 2022; Jo et al., 2022; Aldini et al., 2022; Rashid et al., 2020; Guo et al., 2020; Francis et al., 2021; Abougarair et al., 2021; Penaloza et al., 2018; Nann et al., 2020)
@@
High Temporal Resolution 8 (9.2%) (Baka et al., 2019; Pawuś and Paszkiel, 2022; Aznan et al., 2019; Kim et al., 2021; Farmaki et al., 2022; Li et al., 2019; Zhao et al., 2020b; Ehrlich and Cheng, 2019)
@@
Seamless Feedback 42 (48.3%) (Araujo et al., 2021; Chen et al., 2021; Roy and Bhaumik, 2022; Kim et al., 2021; Farmaki et al., 2022; Li et al., 2019; Prinsen et al., 2022; Roshdy et al., 2021; Bahman and Shamsollahi, 2019; Braun et al., 2019; Li et al., 2023; Qian et al., 2018; Dissanayake et al., 2022; Chang and Sun, 2021; Aldini et al., 2019; Staffa and Rossi, 2022; Iwane et al., 2019; Kompatsiari et al., 2018; Long et al., 2019; Yoon et al., 2021; Ogino and Mitsukura, 2018; Cervantes et al., 2023; Kar et al., 2021; Chu et al., 2018; Kilmarx et al., 2018; Li et al., 2022a; Lu et al., 2020; Zhang et al., 2021; Du et al., 2021; Sanguantrakul et al., 2020; Hernandez-Carmona and Penaloza, 2019; Chiuzbaian et al., 2019; Ghosh and Orlando, 2019; Wei et al., 2021; Ali et al., 2021; Gordleeva et al., 2020; Kuffuor and Samanta, 2018; Memar and Esfahani, 2018; Richter et al., 2023; Abougarair et al., 2021; Penaloza et al., 2018; Nann et al., 2020)
Computational Efficiency 4 (4.6%) (Lu et al., 2022; Rahul and Sharma, 2019; Araujo et al., 2021; Richter et al., 2023)
@@
Feedback Enhancement Visual Feedback 28 (32.2%) (Baka et al., 2019; Belkacem and Lakas, 2021; Farmaki et al., 2022; Sugiyama et al., 2023; Kompatsiari et al., 2018; Si-Mohammed et al., 2018; Chenga et al., 2023; Kilmarx et al., 2018; Wang et al., 2018c; Cao and Liu, 2018; Abougarair et al., 2021; Du et al., 2021; Magee and Givigi, 2021; Penaloza et al., 2018; Yu et al., 2019; Chiuzbaian et al., 2019; Bahman and Shamsollahi, 2019; Ghosh and Orlando, 2019; Mondini et al., 2020; Rahul and Sharma, 2019; Yuan and Li, 2018; Quiles et al., 2022; Prinsen et al., 2022; Aznan et al., 2019; Hernandez-Carmona and Penaloza, 2019; Zhang et al., 2021; Ehrlich and Cheng, 2019; Alimardani et al., 2022)
Auditory Feedback 3 (3.4%) (Kar et al., 2021; Richter et al., 2023; Wang and Sugaya, 2021)
Multimodal Feedback 7 (8%) (Braun et al., 2019; Li et al., 2023; Chen et al., 2021; Li et al., 2022a; Roy and Bhaumik, 2022; Li et al., 2019; Dissanayake et al., 2022)
@@
Direct Feedback 22 (25.3%) (Prinsen et al., 2022; Aznan et al., 2019; Hernandez-Carmona and Penaloza, 2019; Zhang et al., 2021; Ehrlich and Cheng, 2019; Alimardani et al., 2022; Chhabra et al., 2020; Fang et al., 2022; Memar and Esfahani, 2018; Lu et al., 2022; Pawuś and Paszkiel, 2022; Jo et al., 2022; Aldini et al., 2022; Martínez-Cagigal et al., 2020; Sanguantrakul et al., 2020; Wei et al., 2021; Araujo et al., 2021; Francis et al., 2021; Rashid et al., 2020; Staffa and Rossi, 2022; Cervantes et al., 2023; Kar et al., 2021)
@@
Performance Feedback 18 (20.1%) (Cervantes et al., 2023; Guo et al., 2020; Qian et al., 2018; Wu et al., 2022; Iwane et al., 2019; Memar and Esfahani, 2019; Long et al., 2019; Liu and Jebelli, 2021; Lu et al., 2020; Shao et al., 2020; Boonarchatong and Ketcham, 2023; Kuffuor and Samanta, 2018; Jiang et al., 2018; Chen et al., 2020; Wang et al., 2018b; Ai et al., 2018; Lyu et al., 2022; Aldini et al., 2019)
@@
Neurofeedback 22 (25.3%) (Staffa and Rossi, 2022; Kim et al., 2021; Dissanayake et al., 2022; Chang and Sun, 2021; Toichoa Eyam et al., 2021; Alimardani et al., 2020; Yoon et al., 2021; Tang et al., 2022; Ogino and Mitsukura, 2018; Chu et al., 2018; Roshdy et al., 2021; Ali et al., 2021; Xu et al., 2018; Ak et al., 2022; Zhao et al., 2020b; Gordleeva et al., 2020; Nann et al., 2020; Korovesis et al., 2019; Karunasena et al., 2021; Guo et al., 2020; Lyu et al., 2022; Aldini et al., 2019; Richter et al., 2023)
@@
Artificial Intelligence Methods Traditional Machine Learning (ML) 42 (48.3%) (Araujo et al., 2021; Cao and Liu, 2018; Magee and Givigi, 2021; Chen et al., 2021; Roy and Bhaumik, 2022; Farmaki et al., 2022; Li et al., 2019; Belkacem and Lakas, 2021; Chhabra et al., 2020; Xu et al., 2018; Chen et al., 2020; Wang et al., 2018b; Jo et al., 2022; Aldini et al., 2022; Rashid et al., 2020; Guo et al., 2020; Zhao et al., 2020b; Ehrlich and Cheng, 2019; Francis et al., 2021; Braun et al., 2019; Qian et al., 2018; Chang and Sun, 2021; Iwane et al., 2019; Long et al., 2019; Ogino and Mitsukura, 2018; Kar et al., 2021; Chu et al., 2018; Lu et al., 2020; Zhang et al., 2021; Du et al., 2021; Wei et al., 2021; Gordleeva et al., 2020; Kuffuor and Samanta, 2018; Penaloza et al., 2018; Lyu et al., 2022; Memar and Esfahani, 2019; Si-Mohammed et al., 2018; Shao et al., 2020; Boonarchatong and Ketcham, 2023; Yuan and Li, 2018; Jiang et al., 2018; Quiles et al., 2022)
@@
Deep Learning (DL) 19 (21.8%) (Rahul and Sharma, 2019; Magee and Givigi, 2021; Pawuś and Paszkiel, 2022; Aznan et al., 2019; Li et al., 2019; Wu et al., 2022; Tang et al., 2022; Korovesis et al., 2019; Aldini et al., 2022; Guo et al., 2020; Roshdy et al., 2021; Li et al., 2023; Staffa and Rossi, 2022; Kar et al., 2021; Lu et al., 2020; Sanguantrakul et al., 2020; Ghosh and Orlando, 2019; Fang et al., 2022; Chenga et al., 2023)
@@
Non-ML/DL 7 (8%) (Baka et al., 2019; Wang et al., 2018c; Yoon et al., 2021; Li et al., 2022a; Hernandez-Carmona and Penaloza, 2019; Chiuzbaian et al., 2019; Wang and Sugaya, 2021)
@@ 4.3.1. Application Contexts
    Context Category Domain Description Papers (%) References
    Human Control Concern (HCC) Human-centric Technology Usability and user experience Exploring better UX for effective interactions 7 (8%) (Fang et al., 2022; Kim et al., 2021; Martínez-Cagigal et al., 2020; Kuffuor and Samanta, 2018; Yuan and Li, 2018; Memar and Esfahani, 2019; Iwane et al., 2019)
    Research and innovation Exploring new advancements for research needs 18 (20.7%) (Ehrlich and Cheng, 2019; Sugiyama et al., 2023; Kilmarx et al., 2018; Cao and Liu, 2018; Chen et al., 2021; Liu and Jebelli, 2021; Chen et al., 2020; Zhao et al., 2020b; Hernandez-Carmona and Penaloza, 2019; Ak et al., 2022; Guo et al., 2020; Yuan and Li, 2018; Jiang et al., 2018; Si-Mohammed et al., 2018; Pawuś and Paszkiel, 2022; Wei et al., 2021; Cervantes et al., 2023; Lu et al., 2020)
    Mutual collaboration Collaborating with robots to enhance task performance by control 3 (3.4%) (Lyu et al., 2022; Liu and Jebelli, 2021; Memar and Esfahani, 2018)
    Assistance Physical movement Aiding people with motor disorders for physical movements 19 (21.8%) (Farmaki et al., 2022; Zhang et al., 2021; Du et al., 2021; Abougarair et al., 2021; Magee and Givigi, 2021; Wang et al., 2018b; Li et al., 2019; Chiuzbaian et al., 2019; Ghosh and Orlando, 2019; Mondini et al., 2020; Rahul and Sharma, 2019; Ali et al., 2021; Francis et al., 2021; Bahman and Shamsollahi, 2019; Li et al., 2023; Long et al., 2019; Wang et al., 2018c; Tang et al., 2022; Wei et al., 2021)
    Communication Motionless control from people for intended actions 7 (8%) (Boonarchatong and Ketcham, 2023; Shao et al., 2020; Rashid et al., 2020; Korovesis et al., 2019; Karunasena et al., 2021; Quiles et al., 2022; Bahman and Shamsollahi, 2019)
    Healthcare Rehabilitation Rehabilitation for people with mobility impairment 15 (17.2%) (Wu et al., 2022; Chu et al., 2018; Ai et al., 2018; Li et al., 2022a; Roy and Bhaumik, 2022; Jo et al., 2022; Nann et al., 2020; Xu et al., 2018; Sanguantrakul et al., 2020; Araujo et al., 2021; Gordleeva et al., 2020; Chhabra et al., 2020; Dissanayake et al., 2022; Wang et al., 2018c; Tang et al., 2022)
    Therapy Therapeutic process 1 (1.1%) (Roshdy et al., 2021)
    Entertainment Engagement in entertaining activities 1 (1.1%) (Cervantes et al., 2023)
    Safety and Security Exploration for safe and secure needs 1 (1.1%) (Lu et al., 2020)
    Social Interaction Robots as companions for social communications 1 (1.1%) (Toichoa Eyam et al., 2021)
    Military Usage Exploration intended for military cases 1 (1.1%) (Belkacem and Lakas, 2021)
    Robot Design Concern (RDC) Human-centric Technology Usability and user experience Exploring better UX for effective interactions 2 (2.3%) (Aldini et al., 2019; Chenga et al., 2023)
    Research and innovation Exploring new advancements for research needs 5 (5.7%) (Richter et al., 2023; Penaloza et al., 2018; Lu et al., 2022; Kompatsiari et al., 2018; Ogino and Mitsukura, 2018)
    Mutual collaboration Collaborating with robots to enhance task performance 2 (2.3%) (Aldini et al., 2022, 2019)
    Assistance Physical movement Aiding people with motor disorders for physical movements 3 (3.4%) (Yu et al., 2019; Aznan et al., 2019; Yoon et al., 2021)
    Communication Aiding motionless people for intended actions 2 (2.3%) (Braun et al., 2019; Wang and Sugaya, 2021)
    Healthcare Rehabilitation Rehabilitation for people with mobility impairment 1 (1.1%) (Chenga et al., 2023)
    Mental recovery Supports mental recovery 1 (1.1%) (Yoon et al., 2021)
    Meditation Help with mindfulness and meditation process 1 (1.1%) (Yoon et al., 2021)
    Education Language learning Supports exclusively in language learning 2 (2.3%) (Prinsen et al., 2022; Alimardani et al., 2022)
    Self-learning Support self-learning for skill enrichment 3 (3.4%) (Wang and Sugaya, 2021; Kar et al., 2021; Alimardani et al., 2020)
    Entertainment Engagement in entertaining activities 1 (1.1%) (Kar et al., 2021)
    Safety and Security Exploration for safe and secure needs 1 (1.1%) (Qian et al., 2018)
    Social Interaction Robots as companions for social communications 4 (4.6%) (Alimardani et al., 2020; Baka et al., 2019; Chang and Sun, 2021; Staffa and Rossi, 2022)